# Llama-3.1-Sherkala-8B-Chat: An Open Large Language Model for Kazakh

### v1

@@ Abstract
Llama-3.1-Sherkala-8B-Chat, or Sherkala-Chat (8B) for short, is a state-of-the-art instruction-tuned open generative large language model (LLM) designed for Kazakh. Sherkala-Chat (8B) aims to enhance the inclusivity of LLM advancements for Kazakh speakers. Adapted from the LLaMA-3.1-8B model, Sherkala-Chat (8B) is trained on 45.3B tokens across Kazakh, English, Russian, and Turkish. With 8 billion parameters, it demonstrates strong knowledge and reasoning abilities in Kazakh, significantly outperforming existing open Kazakh and multilingual models of similar scale while achieving competitive performance in English. We release Sherkala-Chat (8B) as an open-weight instruction-tuned model and provide a detailed overview of its training, fine-tuning, safety alignment, and evaluation, aiming to advance research and support diverse real-world applications. [^1] Sherkala-Chat (8B) can be accessed at https://huggingface.co/inceptionai/Llama-3.1-Sherkala-8B-Chat. This paper contains examples that may be offensive or triggering to some audiences.
@@ 1 Introduction
The Republic of Kazakhstan, with a population of approximately 20 million, is characterized by its linguistic diversity. The majority of the population (73%) speaks Kazakh, followed by 15% who speak Russian, alongside several other Turkic languages. [^2] https://glottolog.org/ Despite this multilingual landscape, the Kazakh language and culture remain significantly underrepresented in natural language processing (NLP) research and resources [30]. This underrepresentation highlights the need for targeted efforts to develop NLP tools and datasets that better reflect the linguistic and cultural richness of the Kazakh-speaking community.
In this technical report, we aim to address the limited progress of large language models within the context of Kazakhstan by introducing state-of-the-art Kazakh language models, namely Sherkala (8B) and Sherkala-Chat (8B). These transformer-based models are pretrained on 45.3 billions of tokens in Kazakh, Russian, and English, with a primary emphasis on the Kazakh language. While recent advancements in language models have thoughtfully considered multilinguality and demonstrated cross-lingual transfer capabilities, they remain predominantly English-centric (e.g., Falcon [1], PALM [19], the latest Aya [77] and Llama-3.1 [21]). This limitation makes them suboptimal for the Kazakhstan context, as we demonstrate in this report, particularly in knowledge-intensive and reasoning evaluations.
Like many other languages with limited resources [30], the scarcity of high-quality Kazakh data poses a significant challenge in developing Sherkala (8B) and Sherkala-Chat (8B). This limitation affects both the availability of unlabeled text for base-model pretraining and the construction of instruction datasets for the chat model. To address this, we carefully collect as much Kazakh data as possible from various sources (Section 2), in addition to supplementing existing texts with a curated dataset generated through a translation pipeline. We also apply rigorous data filtering and cleaning to ensure high-quality training data. To improve Kazakh representation, we leverage cross-lingual transfer from high-resource languages, including English and Russian. In total, we utilize 45.3 billion tokens, comprising 19.45 billion Kazakh tokens, 19.45 billion English tokens, and 6.4 billion Russian and Turkish tokens. For chat model development, we incorporate datasets from Kazakh, English, and Russian sources.
As part of our commitment to responsible AI development, we include a dedicated safety pretraining stage to mitigate potential harms and align the model with ethical and societal considerations relevant to Kazakhstan. This involves filtering harmful or biased content, reinforcing appropriate responses to sensitive topics, and ensuring compliance with ethical guidelines. By integrating safety measures at the training level, we aim to make Sherkala (8B) a more reliable and suitable tool for real-world applications.
We built Sherkala (8B) based on the Llama-3.1 architecture [21], which follows the standard dense transformer framework [79] while integrating recent advancements such as RoPE positional encoding [69] and grouped-query attention [5]. Additionally, we developed a custom tokenizer designed to balance Kazakh and English, optimizing bilingual text processing. To better accommodate the linguistic characteristics of Kazakh and improve model performance in the Kazakhstan context, we extended the vocabulary by 25%, ensuring broader coverage of Kazakh-specific words and phrases.
To assess Sherkala (8B)’s performance, we conducted evaluations across multiple NLP benchmarks in Kazakh, Russian, and English. For multiple-choice question answering, we focused on reasoning, knowledge retention, misinformation detection, and bias analysis. For the evaluation of text generation, we leveraged internal data sets that cover Kazakhstan’s cultural and public government data, ensuring high relevance to the local context. In addition, we conducted a dedicated safety evaluation using a Kazakh-specific safety dataset we created to assess the model’s ability to handle sensitive topics responsibly. The results indicate that Sherkala-Chat (8B) outperforms existing models of comparable size in Kazakh while maintaining strong competitiveness in English and Russian.
By releasing Sherkala-Chat (8B) openly, we aim to encourage further research and development, fostering innovation and real-world applications that benefit both Kazakh speakers and the broader global community. However, despite our rigorous safety measures, we acknowledge that the model is not entirely immune to risks and limitations. We strongly advise users to exercise caution and conduct additional safety evaluations before deploying the model. To support this, we provide detailed responsible release guidelines in Section 9.
Figure 1: Example text generation with Sherkala-Chat (8B).
2 Pretraining Data
@@
The Kazakh dataset is constructed from a diverse range of sources, including openly available datasets, targeted collection of high-quality native Kazakh content, and contributions from cultural institutions. To further expand the dataset, we incorporate rigorously filtered publicly available documents, alongside carefully curated high-quality educational content and code, and finally the dataset is enriched with synthetic data. These enhancements significantly increase the volume of high-quality Kazakh tokens. Collectively, these sources create a comprehensive repository spanning academic, cultural, and institutional domains, ensuring a well-rounded and representative knowledge base for the model.
For English, Russian, and Turkish, we use openly available, high-quality curated data sources, covering academic and scientific content, textbooks, and encyclopedia ensuring a well-rounded representation of general and domain-specific knowledge.
The composition of these languages in the training mix is carefully determined based on our preliminary experiments, which assess the trade-offs between cross-lingual transfer benefits and potential dilution of Kazakh proficiency. The rationale behind our language mixture strategy, along with its impact on model performance, is discussed in Section 2.2.
2.1 Preprocessing Pipeline
Standardization
Filtering
Cleaning
Re-Filtering
Deduplication
Figure 2: Our Kazakh preprocessing pipeline.
Developing the preprocessing pipeline for Kazakh posed greater challenges compared to English. While English preprocessing pipelines benefit from numerous large-scale, open-access datasets and well-established techniques, Kazakh requires a custom-built approach. Insights gained from experiments with smaller LLMs and the preprocessing pipeline for the dataset used for Jais [72] guided the selection of heuristics used in the final pipeline for Sherkala (8B)’s dataset. Due to the limited availability of Kazakh data, we applied less aggressive filtering than typically used for English, ensuring that valuable Kazakh content was retained.
2.2 Mixing Multilingual Data
During the adaptation of the Llama-3.1-8B model, we primarily mix Kazakh and English data following the findings of [27], while adding Russian and Turkish to maintain practical relevance of the model and enable transfer of capabilities via cross-lingual transfer. Domain adaptation involves continual pre-training of a foundation model on new data not seen during the pre-training. When this new domain data is out-of-distribution for the model being adapted, it can cause significant forgetting of prior capabilities, which is referred to as a stability gap [24]. Adding a small proportion of replay data, which is closer in distribution to the pre-training data, can mitigate forgetting. While the exact composition of pre-training data for Llama-3.1-8B is unknown, we run a set of smaller experiments for adaptation using data mixes from various high quality sources to form an estimate of which datasets mitigate forgetting of the model’s English capabilities. For the final replay data used in Sherkala (8B), we include a mix of synthetic and non-synthetic datasets, primarily focusing on high quality content in Machine Learning, Mathematics, Coding, and academic content in other disciplines.
To determine the optimal proportion in which data in different languages are to be mixed, we performed a series of ablation studies varying the mixed proportions of Kazakh against English (with a smaller proportion of Russian and Turkish). The trained models were evaluated for best Kazakh language performance while maintaining performance in English and Russian. Our findings indicate that maintaining equal proportions of Kazakh and English yields the best results, allowing for improved Kazakh capabilities while preserving the model’s original performance. Reducing the amount of English data led to a noticeable decline in English performance without providing significant gains in Kazakh.
Based on these insights, we adopted a 3:1:3 ratio for Kazakh:Russian+Turkish:English in the final pretraining setup. Llama-3.1-8B was continually pretrained on 45.3 billion tokens, with 19.45 billion tokens each in Kazakh and English, and 6.4 billion tokens in Russian and Turkish combined.
3 Model
3.1 Model Architecture
@@
Sherkala (8B) follows a standard transformer-based architecture [79], adapting and continuing pretraining on the base variant of the Llama-3.1-8B model, a causal decoder-only design similar to those used in GPT-2 [62] and Llama [75]. We selected this model due to Llama-3’s strong multilingual performance across various benchmarks. Our preliminary analysis during Sherkala (8B)’s development indicated that Llama-3.1-8B serves as a better base model for Kazakh pretraining than Llama-3-8B. To extend Llama-3.1 into a Kazakh-specific model, we incorporate several recent advancements from the literature along with optimizations based on our experiments.
Sherkala (8B) Tokenizer:
Although Llama-3.1 is pretrained on multilingual texts, its training data distribution is imbalanced and predominantly skewed towards English. Consequently, the default Llama-3.1 tokenizer is suboptimal for Kazakh and requires modification before continuous pretraining for language adaptation. Although our primary focus is on adapting the model for Kazakh, our training corpus also includes significant amounts of Russian, Turkish, and English data, ensuring that the resulting model remains capable in these languages.
To develop a balanced tokenizer for English, Kazakh, Russian, and Turkish, we first trained monolingual tokenizers for Kazakh, Russian, and Turkish using the byte pair encoding (BPE) method [68]. We then extended the Llama-3.1 vocabulary by incorporating the most frequent tokens extracted from these monolingual tokenizers, ensuring that none of the new tokens overlapped with the original vocabulary. This expansion increases the vocabulary size by 25% compared to the original tokenizer.
[…]
Table 1 compares the vanilla Llama-3.1 tokenizer with our modified version, showing that the original tokenizer requires up to 2.3×2.3$\times$2.3 × more tokens to represent the same Kazakh text. Extending the Llama-3 vocabulary by 25% reduces the fertility score for Kazakh by 56.87% compared to the base tokenizer, resulting in a final score of 2.04. This effectively halves the number of tokens required to represent Kazakh text, improving computational efficiency during training and inference.
@@
Table 1: Intrinsic evaluation of tokenizer performance across different vocabulary sizes. Expanding the tokenizer with additional Kazakh tokens reduces fertility across all three languages compared to the original Llama-3.1 tokenizer, with the most significant reduction observed in Kazakh.
Sherkala (8B) Embedding:
Following the embedding initialization approach in [27], we employ a semantic similarity-based method using Wechsel multilingual initialization [47], where pretrained embeddings such as FastText or OpenAI embeddings are utilized.
@@
For each new Kazakh token added to the Llama-3.1 base vocabulary, we identify the top-$k$ most similar tokens in the base vocabulary based on cosine similarity using embeddings from a pre-trained embedding model. We use OpenAI’s text-embedding-3-large embeddings [32] due to its high quality and strong multilingual capabilities. To initialize the embeddings of the new Kazakh tokens, we take a weighted average of the top-$k$ similar tokens’ base embeddings. After experimenting with different values for the $k$, we achieve the best results with $k=5$. This initialization method is applied to both the embedding and output layers of Sherkala (8B).
3.2 Model and Training Hyperparameters
For the continual pre-training dataset, we pre-process the documents from the source list described in Section 2. This involves tokenizing and packing the documents into sequences with a context length of 8192 tokens. Packing allows multiple documents to fit within a single sequence of the specified length. To indicate document boundaries, we insert the end-of-sequence (EOS) token <|end_of_text|>, enabling the language model to distinguish between unrelated tokens. As in standard pretraining, attention is applied across the entire sequence.
Model Learning Rate Batch Size Warm Up
Sherkala (8B) $1.5e{-4}$ $64e$ 6 $1\%$
Table 2: Training hyperparameter values: the learning rate, batch size, and warm-up ratio
We train Sherkala (8B) using the AdamW optimizer [34] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-5$, and weight decay of 0.1. We scale the gradient norms using a maximum norm clipping value of 1.0. The learning rate schedule comprises a linear warm-up from 0 to the maximum learning rate in 110 steps, followed by a 10$\times$ cosine decay until 11,433 steps. We use a global batch size of 4M tokens.
3.3 Training Infrastructure
The CG-2 hardware and software stack enables training extremely large models using data parallelism by relying on a special execution mode available with Cerebras Wafer Scale Clusters, called weight streaming. Weight streaming fully bypasses the complexity of 3D parallelism on traditional GPU clusters and provides simpler and higher performance scaling.
4 Instruction-Tuning
@@
While autoregressive pretraining on large-scale unlabeled text yields strong performance, instruction fine-tuning is widely employed to enhance generalization across diverse tasks [55]. This process enables the model to better understand and execute user instructions, improving its adaptability to real-world applications. Instruction tuning refines the model’s ability to generate contextually appropriate responses, follow complex multi-turn interactions, and align outputs with user intent. Moreover, it plays a crucial role in enhancing the model’s safety by reducing harmful, biased, or inappropriate outputs. Given these benefits, we apply instruction fine-tuning to Sherkala (8B) to ensure it can effectively handle a range of NLP tasks in both Kazakh and English while maintaining reliability and user alignment. The result of this instruction fine-tuning is Sherkala-Chat (8B).
4.1 Instruction-Tuning Data
To enhance Sherkala (8B)’s applicability in Kazakhstan, we incorporate instruction-tuning data in Kazakh, Russian, and English. While Sherkala (8B) is designed as a Kazakh-centric model, maintaining its proficiency in English and Russian is essential for broader usability. To achieve this, we construct a multilingual instruction-tuning dataset that balances linguistic diversity while prioritizing Kazakh. The dataset consists of approximately 3.5 million prompt-response pairs in Kazakh, 3.8 million in English, and 263 thousand in Russian. Details of the data sources, preprocessing steps, and composition for each language are provided in the following subsections.
4.1.1 General Instruction-Tuning Data
Kazakh
Kazakh, as a low-resource language, has limited high-quality instruction-tuning datasets, and previous efforts have addressed this limitation by using machine translation to adapt English instruction-tuning data. To develop a Kazakh instruction-tuning dataset, we employ two strategies: (1) translating existing English instruction-tuning datasets into Kazakh using machine translation and (2) constructing an in-house dataset from cultural wiki sources and public governmental data. The latter is generated through fact extraction and question generation, followed by human validation to ensure accuracy and relevance [42]. In total, we curate 3.5 million instruction examples in Kazakh, providing a diverse and high-quality resource for instruction fine-tuning. Further details on the data sources are provided in Table 3.
@@ 4.1.1 General Instruction-Tuning Data
English JAIS finetuning [72] Kazakh Yes 34% 2,759,449
ORCA [45] Kazakh Yes 26% 772,269
English JAIS finetuning [72] English No 10.5% 3,354,983
Tulu V3 [40] English No 19.5% 450,817
Grandmaster-PRO-MAX [51] Russian No 5% 262,755
Safety Dataset Kazakh, English Yes, No 95% 99,981
@@
English

### v2

# Sherkala-Chat: Building a State-of-the-Art LLM for Kazakh in a Moderately Resourced Setting

@@ Abstract
Llama-3.1-Sherkala-8B-Chat, or Sherkala-Chat (8B) for short, is a state-of-the-art instruction-tuned open generative large language model (LLM) designed for Kazakh. Sherkala-Chat (8B) aims to enhance the inclusivity of LLM advancements for Kazakh speakers. Adapted from the LLaMA-3.1-8B model, Sherkala-Chat (8B) is trained on 45.3B tokens across Kazakh, English, Russian, and Turkish. With 8 billion parameters, it demonstrates strong knowledge and reasoning abilities in Kazakh, significantly outperforming existing open Kazakh and multilingual models of similar scale while achieving competitive performance in English. To ensure effective and responsible alignment, we leverage translated instruction datasets, a Kazakhstan-specific instruction dataset that is automatically constructed and manually verified, and Kazakh-specific safety data. We release Sherkala-Chat (8B) as an open-weight model, along with a detailed description of its training, alignment, and evaluation, to support research and real-world applications for Kazakh speakers. [^1] Sherkala-Chat (8B) can be accessed at https://huggingface.co/inceptionai/Llama-3.1-Sherkala-8B-Chat.
@@ 1 Introduction
The Republic of Kazakhstan, with a population of approximately 20 million, is a linguistically diverse nation. Kazakh is the state language and spoken by around 80% of the population. Russian remains widely used, with over 80% of Kazakhstanis speaking this language. Additionally, the country hosts several other Turkic and minority languages. [^2] https://en.wikipedia.org/wiki/Languages_of_Kazakhstan, https://glottolog.org/This paper contains examples that may be offensive or triggering to some audiences. Despite Kazakh’s state status and its growing use in education and government spaces, alongside with diminishing presence of Russian as the ethnic Russian population continues to decline, the Kazakh language and culture remain significantly underrepresented in natural language processing (NLP) research and resources (Joshi et al., 2020). This gap highlights the need for developing Kazakh-centric NLP tools and datasets to support the linguistic and cultural identity of the Kazakh-speaking community and empower millions of Kazakh speakers to fully engage with AI-driven technologies in their own language.
In this paper, we introduce two state-of-the-art Kazakh language models—Sherkala (8B) and its instruction-tuned variant Sherkala-Chat (8B)—designed to advance large language model (LLM) capabilities in the underrepresented Kazakh language. These models are derived from LLaMA-3.1 8B model (Dubey et al., 2024) by continuously training on 45.3 billion tokens spanning Kazakh, Russian, and English, with a strong emphasis on Kazakh. Although recent multilingual models such as Falcon (Almazrouei et al., 2023), PaLM (Chowdhery et al., 2023), Aya (Ustun et al., 2024), and LLaMA-3.1 (Dubey et al., 2024) demonstrate notable cross-lingual transfer capabilities, they remain primarily English-centric and often underperform in lower-resourced contexts. As we show in Section 4, these limitations are especially evident in knowledge-intensive and reasoning tasks relevant to Kazakhstan.
As with many low-resource languages (Joshi et al., 2020), the development of Sherkala (8B) and Sherkala-Chat (8B) faces challenges due to the limited availability of high-quality Kazakh data, impacting both pretraining of the base model and instruction tuning. To overcome this, we adopt a two-pronged strategy which combines model adaptation with high-quality data cuartion. On the modeling side, we continuously pretrain Sherkala (8B) starting from the LLaMA-3.1 model(Dubey et al., 2024), leveraging its capacity for cross-lingual transfer. To better support Kazakh, we design a custom tokenizer that balances Kazakh and English, expanding the vocabulary by 25% to capture Kazakh-specific words and expressions more effectively. We also integrate RoPE positional encoding (Su et al., 2021) and grouped-query attention (Ainslie et al., 2023) to enhance model performance. On the data side, we compile Kazakh data from a wide range of sources (Section 2.1) and augment it with a curated dataset generated through a translation pipeline. To ensure quality, all data undergo rigorous filtering and cleaning. The final pretraining corpus comprises 45.3 billion tokens, including 19.45B Kazakh tokens, 19.45B English tokens, and 6.4B Russian and Turkish tokens.
For instruction tuning (Section 3.1), we use multilingual datasets in Kazakh, English, and Russian, incorporating translated instructions and a Kazakhstan-specific instruction dataset that is semi-automatically constructed using GPT-4o and manually verified for relevance and accuracy. To promote responsible AI development, we include a dedicated safety alignment stage (Section 3.2) focused on minimizing potential harms and aligning the model with ethical and cultural norms specific to Kazakhstan. This includes filtering harmful or biased content, reinforcing appropriate responses to sensitive topics, and ensuring the model complies with established ethical guidelines.
Our overall contributions are as follows:
  • We release Sherkala-Chat (8B) as an open, state-of-the-art language model for Kazakh, contributing to the inclusivity of AI technologies in the Kazakh context.
  • We conduct comprehensive evaluations across multiple NLP benchmarks in Kazakh, Russian, and English, demonstrating strong performance in Kazakh while remaining competitive in English and Russian. For multiple-choice question answering, we focus on reasoning, knowledge retention, and misinformation detection. For text generation, we employ internally curated datasets that reflect Kazakhstan’s cultural and public government domains to ensure relevance to local use cases.
  • We perform a dedicated safety evaluation using a newly constructed Kazakh-specific safety dataset to assess the model’s responses to sensitive topics. Our results show that Sherkala-Chat (8B) handles such scenarios more responsibly than comparable models.
2 Pretraining
2.1 Pretraining Data
@@
Figure 1: Kazakh MMLU Accuracy by Language Mixture
The 19.45 billion Kazakh tokens are sourced from a diverse range of texts, including openly available datasets such as Wikipedia, CommonCrawl, CulturaX (Nguyen et al., 2024), and Kazakh news articles. These sources account for 76% of the total Kazakh corpus. To further enrich the dataset, the remaining portion is supplemented with synthetic data generated by translating English Wikipedia articles into Kazakh using Google Translate. Please refer to Appendix B for further details on our pre-processing steps to obtain high-quality Kazakh texts.
For English, Russian, and Turkish, we source textual data primarily from The Pile (Gao et al., 2020) and CommonCrawl, two extensive and diverse datasets widely used in large-scale language model training. The Pile consists of high-quality curated sources, including academic papers, books, Wikipedia, and web content, ensuring a well-rounded representation of general and domain-specific knowledge. CommonCrawl, on the other hand, provides a large-scale snapshot of web content, capturing a broad range of topics, linguistic styles, and real-world text distributions.
Data Mixture
While the exact pretraining data composition of LLaMA-3.1-8B is not publicly available, we conduct a series of smaller-scale adaptation experiments using data mixtures totaling 10B–25B tokens, with a focus on Kazakh performance. Figure 1 presents Kazakh MMLU accuracy across different language mixtures, where 0:0:0 indicates the performance of the original LLaMA-3.1-8B model without further adaptation. We observe that incorporating Kazakh data alone (1:0:0 for Kazakh:Russian+Turkish:English) significantly improves accuracy. However, adding Russian and Turkish reduces performance. Introducing English alongside Kazakh, we find that the 3:1:3 mixture (Kazakh:Russian+Turkish:English) yields the best overall results.
2.2 Continual Pre-training
@@
Sherkala (8B) follows a standard transformer-based architecture (Vaswani et al., 2017), adapting and continuing pretraining on the base variant of the Llama-3.1-8B model (Touvron et al., 2023). We selected this model due to Llama-3.1’s strong multilingual performance across various benchmarks. Our preliminary analysis during Sherkala (8B)’s development indicated that Llama-3.1-8B serves as a better base model for Kazakh pretraining than Llama-3-8B and Llama-2-7B.
Tokenizer
Although LLaMA-3.1 is pretrained on multilingual data, its training distribution is heavily imbalanced and predominantly skewed toward English. As a result, the default LLaMA-3.1 tokenizer is suboptimal for Kazakh and requires modification prior to continual pretraining for effective language adaptation. While our primary goal is to adapt the model for Kazakh, we also consider English, Russian, and Turkish, given their widespread use in Kazakhstan. To create a more balanced tokenizer for these four languages, we first trained separate monolingual tokenizers for Kazakh, Russian, and Turkish using the byte pair encoding (BPE) method Sennrich et al. (2016). We then extended the LLaMA-3.1 vocabulary by incorporating the most frequent tokens from each monolingual tokenizer, ensuring that the added tokens did not overlap with the original vocabulary. This expansion increased the tokenizer’s vocabulary size by 25% relative to the original.
@@
Table 1: Intrinsic evaluation of tokenizer performance across vocabulary sizes. Adding Kazakh tokens reduces fertility in all three languages, with the largest drop in Kazakh.
[…]
Embedding Initialization
@@
Following the embedding initialization approach in Gosal et al. (2024), we employ a semantic similarity-based method using Wechsel multilingual initialization (Minixhofer et al., 2022), where pretrained OpenAI embeddings are utilized. For each new Kazakh token added to the Llama-3.1 base vocabulary, we identify the top-$k$ most similar tokens in the base vocabulary based on cosine similarity using embeddings from a pre-trained embedding model. We use OpenAI’s text-embedding-3-large embeddings (Kusupati et al., 2022) due to its high quality and strong multilingual capabilities. To initialize the embeddings of the new Kazakh tokens, we take a weighted average of the top-$k$ similar tokens’ base embeddings. After experimenting with different values for the $k$, we achieve the best results with $k=5$. This initialization method is applied to both the embedding and output layers of Sherkala (8B).
Training Details
[…]
3 Instruction Tuning
@@
While autoregressive pretraining on large-scale unlabeled text yields strong performance, instruction fine-tuning is widely employed to enhance generalization across diverse tasks (Ouyang et al., 2022). This process enables the model to better understand and execute user instructions, improving its adaptability to real-world applications. Instruction tuning refines the model’s ability to generate contextually appropriate responses, follow complex multi-turn interactions, and align outputs with user intent. Moreover, it plays a crucial role in enhancing the model’s safety by reducing harmful, biased, or inappropriate outputs. Given these benefits, we apply instruction fine-tuning to Sherkala (8B) to ensure it can effectively handle a range of NLP tasks in both Kazakh and English while maintaining reliability and user alignment. The result of this instruction fine-tuning is Sherkala-Chat (8B).
@@
English JAIS finetuning (Sengupta et al., 2023) Kazakh Yes 35.8% 2,759,449
ORCA (Mitra et al., 2024) Kazakh Yes 10.0% 772,269
CultSet and GovSet (Laiyk et al., 2025) Kazakh No 0.2% 10,500
English JAIS finetuning (Sengupta et al., 2023) English No 43.5% 3,354,983
Tulu V3 (Lambert et al., 2024) English No 5.8% 450,817
Grandmaster-PRO-MAX (Nikolich et al., 2024) Russian No 3.4% 262,755
Safety Dataset Kazakh, English Yes, No 1.3% 99,981
@@
3.1 General Instruction Tuning Dataset
To enhance Sherkala (8B)’s applicability in Kazakhstan, we incorporate instruction-tuning data in Kazakh, Russian, and English. While Sherkala (8B) is designed as a Kazakh-centric model, maintaining its proficiency in English and Russian is essential for broader usability. To achieve this, we construct a multilingual instruction-tuning dataset that balances linguistic diversity while prioritizing Kazakh. The dataset consists of approximately 3.5 million prompt-response pairs in Kazakh, 3.8 million in English, and 263 thousand in Russian.
Kazakh:
To develop a Kazakh instruction-tuning dataset, we employ two strategies: (1) translating existing English instruction-tuning datasets into Kazakh using machine translation and (2) constructing an in-house dataset from cultural wiki sources and public governmental data. The latter is generated through fact extraction and question generation, followed by human validation to ensure accuracy and relevance (Laiyk et al., 2025). In total, we curate 3.5 million instruction examples in Kazakh, providing a diverse and high-quality resource for instruction fine-tuning. Further details on the data sources are provided in Table 2.
English: